mathslover (mathslover):

QUIZ TIME : RULES : The questions will be not hard, they all will be easy. The one who answers first and picks it right gets +4 marks, the second one should NOT give answer directly, he must write a little hint for the solution and if it is correct then he will get +2 marks. NEGATIVE MARKING = (-1) So, friends are you ready?

mathslover (mathslover):

Any questions?

terenzreignz (terenzreignz):

I marvel at what you find easy.

terenzreignz (terenzreignz):

But other than that, no

mathslover (mathslover):

:) Also, only 2 users's answers will be considered. Suppose, first answer given is wrong then the user will get (-1) and still there will be 2 chances left for all users to answer. EACH USER CAN ANSWER 2 TIMES (IF HIS/HER FIRST TRY WAS WRONG)

OpenStudy (anonymous):

Ready. Start!

OpenStudy (anonymous):

Ready!

OpenStudy (anonymous):

Who here loves my rating

mathslover (mathslover):

Wait for 1 minute more, so that all come here.

OpenStudy (anonymous):

what.

thomaster (thomaster):

\(\Huge\ddot\smile\)

OpenStudy (anonymous):

Wait, math? *hides under a rock*

terenzreignz (terenzreignz):

satellite's here, oh well, game over... XD

OpenStudy (jhannybean):

Ikr..

terenzreignz (terenzreignz):

oh noes he left 0.0

mathslover (mathslover):

SO FRIENDS HERE IS THE FIRST QUESTION : \(\textbf{If p+2 , 4p-6 and 3p-2 are the three consecutive terms of an AP } \\ \textbf{then find the value of P}\) Best of Luck Question LEVEL - EASY

terenzreignz (terenzreignz):

p = 3

OpenStudy (anonymous):

Hint: You have to connect terms 2 and 1 with 3 and 2

terenzreignz (terenzreignz):

solutions necessary?

terenzreignz (terenzreignz):

oi, @mathslover ?

mathslover (mathslover):

No terenz not necessary and yes you're right. +4

OpenStudy (jhannybean):

Explain? T_T

mathslover (mathslover):

yes u r right @giolefpatceid , +2 for you.

terenzreignz (terenzreignz):

Add them up, and it should be equal to three times the middle term. solve for p algebraically

mathslover (mathslover):

you have to find common difference, 3rd term - 2nd term = 2nd term - 1st term 3p - 8 = -p + 4 4p - 12=0 4p =12 p=3

mathslover (mathslover):

Nice teren. So r u all ready for next question?

OpenStudy (anonymous):

Ready!

terenzreignz (terenzreignz):

sure... :)

mathslover (mathslover):

@terenzreignz : 4 @giolefpatceid :2 ^ Points yet.

mathslover (mathslover):

SECOND QUESTION (total - 10) , what will be the next term of this series : 1,4,10,22, ... QUESTION LEVEL : THINK AND IT WILL BE EASY

OpenStudy (anonymous):

40

OpenStudy (anonymous):

46

mathslover (mathslover):

-1. One more chance for you @giolefpatceid (pm me your process) good work @FutureMathProfessor

OpenStudy (anonymous):

W00T

mathslover (mathslover):

+4 for @FutureMathProfessor and -1 for @giolefpatceid Next answer should contain hint guys

thomaster (thomaster):

1+3 = 4 4+6=10 10+12=22 22+24=46 :3

terenzreignz (terenzreignz):

<not an answer> my internet died at the most inconvenient time T.T

OpenStudy (anonymous):

DDOS

thomaster (thomaster):

start with 3 and double it each time and add to the previous number?

OpenStudy (jhannybean):

yAY! thomas!!

mathslover (mathslover):

So guys here are points: @terenzreignz : 4 @FutureMathProfessor :4 @giolefpatceid : 3 @thomaster : 2

mathslover (mathslover):

Ready for next question?

OpenStudy (anonymous):

It should be 2-1=1 for me

thomaster (thomaster):

lol

OpenStudy (jhannybean):

@giolefpatceid he's being generous :P

mathslover (mathslover):

THIRD QUESTION : IF A, B AND C ARE THE ANGLES OF A TRIANGLE, AND A = 90 , THEN WHAT WILL BE \(\tan(B+C)\)

OpenStudy (anonymous):

undefined

terenzreignz (terenzreignz):

undefined. B+C is 90

terenzreignz (terenzreignz):

crud...

OpenStudy (anonymous):

gg

mathslover (mathslover):

THIRD QUESTION : IF A, B AND C ARE THE ANGLES OF A TRIANGLE, AND A = 90 , THEN WHAT WILL BE tan(B+C)

thomaster (thomaster):

domain error xD

terenzreignz (terenzreignz):

still crud :D

OpenStudy (anonymous):

undefined

OpenStudy (anonymous):

this is legend. haha

OpenStudy (anonymous):

again

OpenStudy (anonymous):

I was the first one @mathslover

OpenStudy (anonymous):

looool

OpenStudy (anonymous):

B+C = 90º , so, tan(B+C) = tan90º and it's undefined.

OpenStudy (jhannybean):

lol he rubs it in people's faces x_x

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

I should get 4 points for getting the answer and @terenzreignz should get 2 points for giving the hint/reason

OpenStudy (jhannybean):

Zair showed his work xDD

terenzreignz (terenzreignz):

You guys should play nicer with each other. :P

OpenStudy (anonymous):

"If you head in the air at a 90 degree angle in Los Angeles California, how high will you be when you are directly overhead Las Vegas Nevada" the ultimate tangent problem

terenzreignz (terenzreignz):

Well, where in the blazes is @mathslover ?

OpenStudy (anonymous):

We DDOS'ed his computer LOL

OpenStudy (jhannybean):

you wont be very high considering the amount of pollution....x_x

OpenStudy (anonymous):

So in other words..I'd be passed out I'd be so high from all the low-quality air

OpenStudy (jhannybean):

Yes XD

OpenStudy (jhannybean):

You wouldnt even REACH nevada,let alone be able to pass over Blithe, AZ....

OpenStudy (anonymous):

Hey guys since I got the right answer you all should medal me

terenzreignz (terenzreignz):

For that, I"m not even going to think about it XD

terenzreignz (terenzreignz):

I'm*

mathslover (mathslover):

Good work @FutureMathProfessor +4 and @zairhenrique : good : +2

mathslover (mathslover):

[OS STOPPED WORKING, SO THANKS FOR WAITING]

OpenStudy (anonymous):

ddos

mathslover (mathslover):

POINTS : @FutureMathProfessor (1st place) : 8 @terenzreignz (2nd place ) :4 @giolefpatceid (3rd place ) :3 @thomaster and @zairhenrique (4th place) : 2

mathslover (mathslover):

So ready for next question, friends?

OpenStudy (anonymous):

Yeah kids this is how we do it

OpenStudy (anonymous):

Ready!

OpenStudy (anonymous):

Yep!

terenzreignz (terenzreignz):

0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

One of the variables is always 0 so the matrix must be 0

mathslover (mathslover):

Find the value of the above determinant

OpenStudy (anonymous):

:(

thomaster (thomaster):

lol wut is a matrix XD

terenzreignz (terenzreignz):

matrix, you know, the one with neo and stuff :P

thomaster (thomaster):

*dodges bullets*

OpenStudy (anonymous):

Why did you delete your post? I was solving it

thomaster (thomaster):

the answer is chicken sandwich

mathslover (mathslover):

@terenzreignz : + 4 @zairhenrique : +2 (though u didn't give the hint, but still I will consider that )

terenzreignz (terenzreignz):

no need to solve, if it has an entire row or column of zeros, the determinant must be zero...

OpenStudy (anonymous):

@mathslover I gave a hint though! You saw it!

OpenStudy (anonymous):

@mathslover Scroll up! Those points are mine!